find all values of y which satisfy the equation
[(y+7)/(y+1)] +1
=[(y+7)/(y-1)]

Answer by stanbon(75887)


[(y+7)/(y+1)]-[(y+7)/(y-1)]=-1
Simplify:
[(y^2+6y-7-(y^2+8y+7)]=-[y^2-1]
y^2-2y-15=0
(y-5)(y+3)=0
Y=5 OR Y=-3